If cos(πœƒ ) = 1/3 and tan( πœƒ ) < 0 , find the exact value of each of the remaining trigonometric functions of πœƒ


1
Expert's answer
2021-03-31T15:32:30-0400

if cos "\theta" >0 and tan "\theta" <0 then this falls on the fourth quadrant. The sin will also be negative.

let x be the adjacent, y the opposite and r the hypotenuse of a right angled triangle where "\theta" is the angle formed as x and r intersect.

"cos \theta=\frac x r"

x=1, r=3

"y=\sqrt {r^2-x^2}"

"y=\sqrt{3^2-1^2}=\sqrt 8"

"sec \theta=\frac{1}{cos \theta}=\frac{1}{\frac 1 3}=3"

"\sin \theta =-\frac y r=- \frac {\sqrt{8}}3"

"\csc \theta =\frac 1 {sin\theta}=\frac {1}{-\frac {\sqrt 8}{3}}=-\frac{3}{\sqrt 8}"

"\tan \theta =-\frac y x=-\frac {\sqrt 8} 1=-\sqrt 8"

"\cot \theta=\frac{1}{tan \theta}=-\frac{1}{\sqrt 8}"